OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the points (6,7) and (2,5)?

OpenStudy (anonymous):

1. calculate for the slope: \[slope(m)=\frac{ y _{2} -y _{1}}{ x _{1}-x _{2} }\] 2. using one of the points (x1, y1 and the slope (m), find the equation of the line by plugging them into \[y-y _{1}=m(x-x _{1})\] and simplify it.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

np:)

OpenStudy (calculusxy):

Explaining this method is kind of hard, but I will try to explain it to you the best as I can. The method that I personally use is rise/run. What I do in this is think about how I can connect the point on the y axis to the x axis. First, I count how many steps up/down do I have to do get to the x point using the vertical axis. So in this case you have to go down 4 steps meaning -4 as the rise. To do the run, I count how many steps I have to count to get to the x point. In this case I have to go right in two counts. So rise/run=-4/2 or also -2